An object starts from rest and accelerates at a rate of 2 m/s squared over a distance of 20 m what is the final velocity
1
Expert's answer
2020-05-13T09:48:25-0400
"v^2=2as\to v^2=2(2)(20)"

"v=8.9\frac{m}{s}"